Miller Toy Company manufactures a plastic swimming pool at its Westwood Plant. The plant has been experiencing problems as shRequired: 1. Compute the following variances for June: a. Materials price and quantity variances. b. Labor rate and efficiencRequired: 1. Compute the following variances for June: a. Materials price and quantity variances. b. Labor rate and efficiencNEED HELP ASAP PLEASE
Answer #1
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($2.00/Pound - $2.45/Pound ) × 15800 Pound = $7110 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((3.6 Pound X 3000)-10600Pound )X $2 = $400 F
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($6.6/hour - $6.3/hour) × 2100 Hours= $630F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (0.5 Hour × 3000 - 2100) × $6.6 = $3960 U F
Calculation of Variable OH rate variance
= (Standard Variable OH per hour - actual variable OH per hour) × Actual hours used
= ($2.10/hour - $2.5/hour) × 1200 Hours= $480 U F
* Actual Cost /Hour= $3000/1200=$2.5/Hour
Calculation of Variable OH efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (0.3Hour × 3000 - 1200) × $2.10 = $630 U F
Part-2
Computation of Net variance
Material pprice Variance 7110UF
Materal Quantity Variance 400F
Labour Rate Variance 810 F
Labor Efficiency Variance 3960UF
Variable Overhead Rate Variance 480UF
Variable Overhead Efficiency Variance 630UF
Net Variance 10970 UF